# Information Entropy Invariance: Enhancing Length Extrapolation in Attention Mechanisms

### v1

@@ Abstract
Improving the length extrapolation capabilities of Large Language Models (LLMs) remains a critical challenge in natural language processing. Many recent efforts have focused on modifying the scaled dot-product attention mechanism, and often introduce scaled temperatures without rigorous theoretical justification. To fill this gap, we introduce a novel approach based on information entropy invariance. We propose two new scaled temperatures to enhance length extrapolation. First, a training-free method InfoScale is designed for dot-product attention, and preserves focus on original tokens during length extrapolation by ensuring information entropy remains consistent. Second, we theoretically analyze the impact of scaling (CosScale) on cosine attention. Experimental data demonstrates that combining InfoScale and CosScale achieves state-of-the-art performance on the GAU-α model with a context window extended to 64 times the training length, and outperforms seven existing methods. Our analysis reveals that significantly increasing CosScale approximates windowed attention, and highlights the significance of attention score dilution as a key challenge in long-range context handling. The code and data are available at https://github.com/HT-NEKO/InfoScale.
@@ 1 Introduction
Transformer-based Large Language Models (LLMs) (Vaswani, 2017) have achieved remarkable success in various Natural Language Processing (NLP) tasks, particularly those requiring long-range context understanding. However, the limited context window imposed by the training data remains a significant bottleneck. This limitation arises partly from the dilution of attention scores by newly added tokens, which disrupts the original distribution and hinders effective long-range modeling (Su, 2021).
@@
To further enhance performance, we extend our analysis to cosine attention. By integrating InfoScale with cosine attention, we achieve significant improvements over existing length extrapolation techniques, including RoPE-based methods (PI (Chen et al., 2023b), YaRN (Peng et al., 2023), CLEX (Chen et al., 2023a), ReRoPE (Su, 2023b)), bias methods (ALiBi (Press et al., 2021)), window-based methods (StreamingLLM (Xiao et al., 2023), LM-Infinite (Han et al., 2024), and unpublished Windowed Attention), and skip-wise training methods (PoSE (Zhu et al., 2023)).
Furthermore, we theoretically analyze the behavior of scaled cosine attention (CosScale) and validate these theoretical findings with empirical evidence. We demonstrate that as CosScale increases, the attention mechanism approximates windowed attention. This observation suggests that attention score dilution poses a more significant challenge than insufficient training for extending the context window.
@@ 2.1 Scaled Dot-Product Attention
Where $d_{model}=H\times d_{q}=H\times d_{k}=H\times d_{v}$, $H$ is the number of attention heads.
@@
Where for every vector in $\bm{K}$, $\bm{q}_{i}$, $\bm{k}_{j}$ are considered as column vectors so that $\bm{q}_{i}\bm{k}_{j}^{T}$ is simply the Euclidean inner product. Since softmax will transform the summarization of $\bm{q}_{i}\bm{K}^{T}$ to 1, we rewrite the function as
@@
Where ${\sum_{j=1}^{n}\beta_{ij}}=1$, $\beta_{ij}\in[0,1]$, $\beta_{ij}=\frac{e^{\lambda\bm{q}_{i}\bm{k}_{j}^{T}}}{\sum\limits_{j=1}^{n}e^{\lambda\bm{q}_{i}\bm{k}_{j}^{T}}}$.
@@ 2.3 Scaled Cosine Attention
Where $\bm{S}_{ij}$ is the vector norm product of $\bm{q}_{i}$ and $\bm{k}_{j}$, $\cos\theta_{ij}$ is the cosine of the angle between the two vectors, $\bm{q}_{i}$ and $\bm{k}_{j}$, also known as the cosine similarity.
@@
If we define $\bm{S}_{ij}$ as a constant $\bm{S}$ for any pair of column vectors during training, this is referred to as scaled cosine attention. The choice $\bm{S}$ is delicate. Making it too small can lead to vanishing gradients, while setting it too large may cause overly sharp attention scores. The latter, however, can be advantageous for tasks involving length extrapolation.
@@ 2.4 Related Work
Pre-softmax Scaling: The pre-softmax scaling approach proposed by (Peng et al., 2023) for the YaRN model experimentally showed that, for an appropriate scaling factor t, the model could achieve better perplexity scores across extended context windows. However, this method is tightly integrated with YaRN, which dynamically modifies the frequency of each dimension of RoPE during inference. As our method is independent of RoPE, we exclude it from the comparison.
@@ 4 Experiments
This section evaluates the effectiveness of InfoScale and CosScale in handling long contexts and improving the length extrapolation capabilities of LLMs. We benchmark our proposed methods against a diverse set of baselines, including: RoPE-based methods (PI (Chen et al., 2023b), YaRN (Peng et al., 2023), ReRoPE (Su, 2023b)), Bias methods (ALiBi (Press et al., 2021)), Window-based methods (StreamingLLM (Xiao et al., 2023), LM-Infinite (Han et al., 2024), Windowed Attention), and Skip-wise training method (PoSE (Zhu et al., 2023)). Implementation details for each method are provided in Section A.2. We assess the performance of InfoScale and CosScale using two metrics: perplexity (PPL) and masked-token prediction accuracy (ACC), whose detailed definitions are in Section A.3. Evaluation is conducted on sequences up to 64 times longer than the maximum sequence length encountered during training, allowing us to rigorously assess length extrapolation capabilities.
@@ 4.1.2 Experiment Settings
The training process consists of 400 epochs with a warm-up period comprising 10% of the total training steps. We use a linear learning rate scheduler with an initial learning rate of 2e-5. The maximum sequence length during training is 64, and a global batch size of 128 is distributed across 4 Tesla V100-SXM2-32GB GPUs. Optimization is performed using the AdamW optimizer with default hyperparameters. Details on the model training loss can be found in Section A.2.
@@ 4.2.1 InfoScale Demonstrates Length Extrapolation Capabilities
When evaluated on sequences of length 4k, InfoScale significantly improves the performance of several baseline methods. For instance, InfoScale boosts the accuracy (ACC) of ReRoPE (Su, 2023b) by nearly tenfold while simultaneously halving its perplexity (PPL) (see Table 1 and Table 7). While InfoScale consistently improves ACC for most baselines, LM-Infinite and Windowed Attention (W.A.) exhibit a slight increase in PPL despite improved ACC. This discrepancy may be attributed to the assumption in InfoScale’s derivation that high-dimensional word embeddings are distributed on a fixed-radius hypersphere, an assumption not strictly satisfied by scaled dot-product attention. To address this, we introduce cosine attention, which enforces this constraint on the embedding distribution. (A detailed discussion of the results is provided in Section A.7. )
@@ 4.2.2 CosScale Demonstrates Strong Length Extrapolation Capabilities
It is worth noting that we set the CosScale for windowed attention to 16 (Su, 2023a), rather than 128 as used for other algorithms. This is because the context window for the Windowed Attention is 64, which falls outside the long-range context regime where CosScale excels.
These findings suggest that constraining the distribution of word embeddings to a hypersphere facilitates more effective angular training of the attention matrix. Based on these results, CosScale alone exhibits strong length extrapolation capabilities. Furthermore, since InfoScale assumes word embeddings are distributed on a hypersphere, combining CosScale with InfoScale is expected to yield even greater performance improvements.
@@ 4.2.3 Combining InfoScale and CosScale for Enhanced Performance
Table 1: Perplexity (PPL) and accuracy (ACC) for baseline methods without or with both CosScale and InfoScale incorporated, evaluated with sequence lengths ranging from 128 to 4k on the length extrapolation task. Lower PPL and higher ACC indicate better performance. CosScale is set to 128 for all methods except Windowed Attention, where it is set to 16. W.A. denotes Windowed Attention, while I.S. and C.S. represent InfoScale and CosScale, respectively.
@@ 4.3 Behaviors of CosScale
4.3.1 CosScale Approximates Windowed Attention as It Increases
@@
We observe that the histogram of the query-key multiplication (QK) before applying RoPE becomes progressively sharper with larger CosScale values. This results in the attention matrix adopting a shape similar to that of Windowed Attention, where attention is primarily focused on a local neighborhood around each token. Notably, the peak value of the attention histogram decreases at CosScale = 600, likely due to numerical precision limitations exceeding the model’s capacity.
@@ 4.3.2 Larger CosScale Favors Longer Context Windows
4.4 Comparison of Windowed Attention and Cosine Attention with Large CosScale
@@ A.7 Performance of Published Methods on GAU-α
For window-based methods: Windowed Attention achieves the best performance among the three. StreamingLLM places extra emphasis on initial tokens, which may not always be semantically significant (Xiao et al., 2023). This, combined with the truncation of high-frequency signals by ReRoPE, may contribute to its lower performance. LM-Infinite combines ReRoPE and Λ-shaped attention, with a focus on initial tokens. However, its performance is hindered by the lack of ”attention sink” behavior in our experiments, suggesting that the attention mechanism in GAU-α differs from the models used in (Han et al., 2024).
@@ A.8 Additional Tables and Charts
Table 7: Perplexity (PPL) and accuracy (ACC) for baseline methods and their counterparts with InfoScale incorporated, evaluated on the GAU-α model with sequence lengths ranging from 128 to 4k on the length extrapolation task. For simplicity, W.A. denotes the baseline algorithm Windowed Attention, and I.S. represents our method InfoScale.
@@
Table 8: Perplexity (PPL) and accuracy (ACC) for baseline methods and their counterparts with CosScale incorporated, evaluated with sequence lengths ranging from 128 to 4k on the length extrapolation task. Lower PPL and higher ACC indicate better performance. CosScale is set to 128 for all methods except windowed attention, where it is set to 16. W.A. denotes Windowed Attention, and C.S. represents CosScale.
@@
Figure 7: Heatmaps of attention matrices extrapolated to 4096. Panel (a) shows GAU-α with CosScale = 16 applied to windowed attention, while panel (b) presents GAU-α with CosScale = 128 applied to cosine attention. The x- and y-axes represent token positions. A 20x20 local region (from index 3516 to 3535) is extracted from the left heatmap in panel (a) for detailed visualization. The color bar of the left heatmap is scaled to match the value range of the extracted region.

### v2

@@ Abstract
Since the emergence of research on improving the length extrapolation capabilities of large models in 2021, some studies have made modifications to the scaling factor in the scaled dot-product attention mechanism as part of their proposed methods without rigorous theoretical justifications. To fill this gap, we propose two new scaled temperatures based on entropy invariance to enhance length extrapolation. First, a training-free method InfoScale is designed for dot-product attention, and preserves focus on original tokens during length extrapolation by ensuring consistent entropy. Second, we theoretically analyze the impact of scaling (CosScale) on cosine attention. Experimental data demonstrates that combining InfoScale and CosScale achieves state-of-the-art performance on the GAU-α model with a context window extended to 64 times the training length, and outperforms seven existing methods. Our analysis reveals that significantly increasing CosScale approximates the Windowed Attention, and highlights the significance of attention score dilution as a key challenge in long-range context handling. The code and data are available at https://github.com/HT-NEKO/InfoScale.
@@ 1 Introduction
Transformer-based Large Language Models (LLMs) (Vaswani, 2017) have achieved remarkable success in various Natural Language Processing (NLP) tasks, particularly those requiring long-range context understanding. However, the limited-length context window imposed by the training data remains a significant bottleneck. This limitation arises partly from the dilution of attention scores by newly added tokens, which disrupts the original distribution and hinders effective long-range modeling (Su, 2021).
@@
To further enhance performance, we extend our analysis to cosine attention. By integrating InfoScale with cosine attention, we achieve significant improvements over existing length extrapolation techniques, including RoPE-based methods (PI (Chen et al., 2023b), YaRN (Peng et al., 2023), CLEX (Chen et al., 2023a), ReRoPE (Su, 2023b)), bias methods (ALiBi (Press et al., 2021)), window-based methods (StreamingLLM (Xiao et al., 2023), LM-Infinite (Han et al., 2024), and the Windowed Attention), and skip-wise training methods (PoSE (Zhu et al., 2023)).
Furthermore, we theoretically analyze the behavior of scaled cosine attention (CosScale) and validate these theoretical findings with empirical evidence. We demonstrate that as CosScale increases, the attention mechanism approximates the Windowed Attention. This observation suggests that attention score dilution poses a more significant challenge than insufficient training for extending the context window.
@@ 2.1 Scaled Dot-Product Attention
where $d_{model}=H\times d_{q}=H\times d_{k}=H\times d_{v}$, and $H$ is the number of attention heads.
@@
where for every vector in $\bm{K}$, $\bm{q}_{i}$ and $\bm{k}_{j}$ are considered as column vectors so that $\bm{q}_{i}\bm{k}_{j}^{T}$ is simply the Euclidean inner product. Since softmax will transform the summarization of $\bm{q}_{i}\bm{K}^{T}$ to 1, we rewrite the function as
@@
where ${\sum_{j=1}^{n}\beta_{ij}}=1$, $\beta_{ij}\in[0,1]$ and $\beta_{ij}=\frac{e^{\lambda\bm{q}_{i}\bm{k}_{j}^{T}}}{\sum\limits_{j=1}^{n}e^{\lambda\bm{q}_{i}\bm{k}_{j}^{T}}}$.
@@ 2.3 Scaled Cosine Attention
where $\bm{S}_{ij}$ is the vector norm product of $\bm{q}_{i}$ and $\bm{k}_{j}$, and $\cos\theta_{ij}$ is the cosine of the angle between the two vectors, $\bm{q}_{i}$ and $\bm{k}_{j}$, also known as the cosine similarity.
@@
If we define $\bm{S}_{ij}$ as a constant $\bm{S}$ for any pair of column vectors during training, this is referred to as scaled cosine attention. The choice of $\bm{S}$ is delicate. Making it too small can lead to vanishing gradients, while setting it too large may cause overly sharp attention scores. The latter, however, can be advantageous for tasks involving length extrapolation.
@@ 2.4 Related Work
Pre-softmax Scaling: The pre-softmax scaling approach proposed by (Peng et al., 2023) for the YaRN model experimentally showed that, for an appropriate scaling factor, the model could achieve better perplexity scores across extended context windows. However, this method is tightly integrated with YaRN, which dynamically modifies the frequency of each dimension of RoPE during inference. As our method is independent of RoPE, we exclude it from the comparison.
@@ 4 Experiments
This section evaluates the effectiveness of InfoScale and CosScale in handling long contexts and improving the length extrapolation capabilities of LLMs. We benchmark our proposed methods against a diverse set of baselines, including: RoPE-based methods (PI (Chen et al., 2023b), YaRN (Peng et al., 2023), ReRoPE (Su, 2023b)), Bias methods (ALiBi (Press et al., 2021)), Window-based methods (StreamingLLM (Xiao et al., 2023), LM-Infinite (Han et al., 2024), the Windowed Attention), and Skip-wise training method (PoSE (Zhu et al., 2023)). Implementation details for each method are provided in Section A.2. We assess the performance of InfoScale and CosScale using two metrics: perplexity (PPL) and masked-token prediction accuracy (ACC), whose detailed definitions are in Section A.3. Evaluation is conducted on sequences up to 64 times longer than the maximum sequence length encountered during training, allowing us to rigorously assess length extrapolation capabilities.
@@ 4.1.2 Experiment Settings
The training process consists of 400 epochs with a warm-up period comprising 10% of the total training steps. We use a linear learning rate scheduler with an initial learning rate of 2e-5. The maximum sequence length during training is 64, and a global batch size of 128 is distributed across 4 Tesla V100-SXM2-32GB GPUs. Optimization is performed using the AdamW optimizer with default hyperparameter values. Details on the model training loss can be found in Section A.2.
@@ 4.2.1 InfoScale Demonstrates Length Extrapolation Capabilities
When evaluated on sequences of length 4k, InfoScale significantly improves the performance of several baseline methods. For instance, InfoScale boosts the accuracy (ACC) of ReRoPE (Su, 2023b) by nearly tenfold while simultaneously halving its perplexity (PPL) (see Table 1 and Table 7). While InfoScale consistently improves ACC for most baselines, LM-Infinite and the Windowed Attention (W.A.) exhibit a slight increase in PPL despite improved ACC. This discrepancy may be attributed to the assumption in InfoScale’s derivation that high-dimensional word embeddings are distributed on a fixed-radius hypersphere, an assumption not strictly satisfied by scaled dot-product attention. To address this, we introduce cosine attention, which enforces this constraint on the embedding distribution. (A detailed discussion of the results is provided in Section A.7. )
@@ 4.2.2 CosScale Demonstrates Strong Length Extrapolation Capabilities
It is worth noting that we set the CosScale for the Windowed Attention to 16 (Su, 2023a), rather than 128 as used for other algorithms. This is because the context window for the Windowed Attention is 64, which falls outside the long-range context regime where CosScale excels.
These findings suggest that constraining the distribution of word embeddings to a hypersphere facilitates more effective angular training of the attention matrix. Based on these results, CosScale alone exhibits strong length extrapolation capabilities. Furthermore, since InfoScale assumes that the word embeddings are distributed on a hypersphere, combining CosScale with InfoScale is expected to yield even greater performance improvements.
@@ 4.2.3 Combining InfoScale and CosScale for Enhanced Performance
Table 1: Perplexity (PPL) and accuracy (ACC) for baseline methods without or with both CosScale and InfoScale incorporated, evaluated with sequence lengths ranging from 128 to 4k on the length extrapolation task. Lower PPL and higher ACC indicate better performance. CosScale is set to 128 for all methods except the Windowed Attention, where it is set to 16. W.A. denotes the Windowed Attention, while I.S. and C.S. represent InfoScale and CosScale, respectively.
@@ 4.3 Behaviors of CosScale
4.3.1 CosScale Approximates the Windowed Attention as It Increases
@@
We observe that the histogram of the query-key multiplication ($QK^{T}$) before applying RoPE becomes progressively sharper with larger CosScale values. This results in the attention matrix adopting a shape similar to that of the Windowed Attention, where attention is primarily focused on a local neighborhood around each token. Notably, the peak value of the attention histogram decreases at CosScale = 600, likely due to numerical precision limitations exceeding the model’s capacity.
@@ 4.3.2 Larger CosScale Favors Longer Context Windows
4.4 Comparison of the Windowed Attention and Cosine Attention with Large CosScale
@@ A.7 Performance of Published Methods on GAU-α
For window-based methods: the Windowed Attention achieves the best performance among the three. StreamingLLM places extra emphasis on initial tokens, which may not always be semantically significant (Xiao et al., 2023). This, combined with the truncation of high-frequency signals by ReRoPE, may contribute to its lower performance. LM-Infinite combines ReRoPE and Λ-shaped attention, with a focus on initial tokens. However, its performance is hindered by the lack of ”attention sink” behavior in our experiments, suggesting that the attention mechanism in GAU-α differs from the models used in (Han et al., 2024).
@@ A.8 Additional Tables and Charts
Table 7: Perplexity (PPL) and accuracy (ACC) for baseline methods and their counterparts with InfoScale incorporated, evaluated on the GAU-α model with sequence lengths ranging from 128 to 4k on the length extrapolation task. For simplicity, W.A. denotes the baseline algorithm the Windowed Attention, and I.S. represents our method InfoScale.
@@
Table 8: Perplexity (PPL) and accuracy (ACC) for baseline methods and their counterparts with CosScale incorporated, evaluated with sequence lengths ranging from 128 to 4k on the length extrapolation task. Lower PPL and higher ACC indicate better performance. CosScale is set to 128 for all methods except the Windowed Attention, where it is set to 16. W.A. denotes the Windowed Attention, and C.S. represents CosScale.
@@
Figure 7: Heatmaps of attention matrices extrapolated to 4096. Panel (a) shows GAU-α with CosScale = 16 applied to the Windowed Attention, while panel (b) presents GAU-α with CosScale = 128 applied to cosine attention. The x- and y-axes represent token positions. A 20x20 local region (from index 3516 to 3535) is extracted from the left heatmap in panel (a) for detailed visualization. The color bar of the left heatmap is scaled to match the value range of the extracted region.